I tried doing the first part of this problem, but not sure if it is right. I dont know how to do part b. Can you please help me?
1) Using the quadratic equation x2 – 4x – 5 = 0, perform the following tasks:
a) Solve by factoring.
Answer:x=5, -1
Show work in this space. 1x2-4x-5 = (x-5)(x+1)


b) Solve by using the quadratic formula.
Answer:
Show work in this space

Answer by stanbon(75887)

1) Using the quadratic equation x2 – 4x – 5 = 0, perform the following tasks:
a) Solve by factoring.
Answer:x=5, -1
Show work in this space. 1x2-4x-5 = (x-5)(x+1)

b) Solve by using the quadratic formula.
Answer:
Quadratic formula: x=[-b+-sqrt(b^2-4ac]/2a
-----------
x = [4+-sqrt(4^2-4*1*-5]/2
x = [4+-sqrt(36)]/2
x = [4+6]/2 or x = [4-6]/2
x = 10/2 or x = -2/2
x = 5 or x = -1
------------
Cheers,
Stan H.